# Site-selective renormalization and competing magnetic instabilities in paramagnet ${\mathrm{Y}}_{3}{\mathrm{Cu}}_{2}{\mathrm{Sb}}_{3}{\mathrm{O}}_{14}$

### 18 Mar 2026 v1

@@ II.4 Spin Susceptibility
Figure 3: Fermi surface and spin susceptibility $\chi(\bm{q},i\omega_{m}=0)$. (a). The Fermi surface in 3D BZ and in a 2D cut through $\Gamma$ point displays multiple pockets and nesting structures. (b). The spin susceptibility calculated from Eq. (2) displays multiple peaks with only small variation, which is further suppressed by electronic correlations. The blue and yellow curves correspond to different interaction parameters as shown by the legend. (c). The leading eigenvalues of $\chi_{0}U_{s}$ follow a similar structure as $\chi_{s}$ in (b), whose variation is also reduced by electronic correlations. (d). Both the leading eigenvalues of $\chi_{0}$ (the left-yaxis in red) and $\chi_{0}U_{s}$ (the right-yaxis in blue) for different momentum wave vectors fall into a single curve under the evolution of electronic correlations, with the former decreases and the latter increases with $U$.
The renormalization arising from electronic correlations also influences two-particle excitations. To investigate this effect, we examine the spin susceptibility and its evolution under the influence of Coulomb interactions. The spin susceptibility, which encodes the magnetic response of the system, often serves as a key indicator of spontaneous symmetry breaking. In particular, a divergent spin susceptibility at a specific wave vector in the Brillouin zone (BZ) directly reveals the dominant real-space spin correlation pattern associated with that wave vector. Given the absence of long-range magnetic order down to very low temperatures in ${\mathrm{Y}}_{3}{\mathrm{Cu}}_{2}{\mathrm{Sb}}_{3}{\mathrm{O}}_{14}$, we anticipate a spin susceptibility that remains relatively broad and structureless across the entire BZ.
@@
From the Lehmann representation of Eq. (II.4), it is evident that the Fermi surface topology plays a crucial role in determining the divergent behavior of the spin susceptibility. When a substantial number of electrons on the Fermi surface scatter with the same wave vector, the resulting coherence can lead to a divergence in the spin susceptibility at that wave vector. Figure 3(a) displays the three-dimensional Fermi surface, with a two-dimensional cut through the $\Gamma$ point shown on the right. The Fermi surface of ${\mathrm{Y}}_{3}{\mathrm{Cu}}_{2}{\mathrm{Sb}}_{3}{\mathrm{O}}_{14}$ consists of several pockets of comparable size, giving rise to multiple competing nesting vectors. A few representative nesting vectors are indicated in the two-dimensional plot, each potentially contributing comparable coherence to electron scattering. As a result, the spin susceptibility exhibits a multi-peak structure with several competing maxima.
@@
The growing competition among different wave vectors in $\chi_{s}(\bm{Q})$ is further reflected in the leading eigenvalues of $U_{s}\chi_{0}(\bm{Q})$, which appear in the denominator of Eq. (2) and drive the divergence of $\chi_{s}(\bm{Q})$ as they approach unity. Figure 3(c) shows these leading eigenvalues for the same interaction parameters used in panel (b). The variation of $\lambda_{\text{max}}(U_{s}\chi_{0}(\bm{Q}))$ across different $\bm{Q}$ is about 6% for $U=0.5$ eV, but reduces to 1.25% when $U$ is increased to 1.0 eV, despite the absolute value rising from 0.68 to 0.82. Consistent with the behavior of $\chi_{s}(\bm{Q})$, increasing interaction strength suppresses the momentum dependence of $\lambda_{\text{max}}(U_{s}\chi_{0}(\bm{Q}))$, leading to a nearly simultaneous divergence tendency at all $\bm{Q}$.
Figure 3(d) extends this analysis by displaying the leading eigenvalues for several values of $U$. In addition to $U_{s}\chi_{0}(\bm{Q})$ (right y-axis, blue), we also show the leading eigenvalues of $\chi_{0}(\bm{Q})$ (see Eq. (II.4) and the left y-axis of Fig. 3(d) in red), which appear in the numerator of Eq. (2). As $U$ increases, the leading eigenvalues of $U_{s}\chi_{0}(\bm{Q})$ at characteristic wave vectors grow with nearly identical magnitudes. Concurrently, those of the bubble susceptibility $\chi_{0}(\bm{Q})$ decrease. Notably, for both quantities, the leading eigenvalues across all $\bm{Q}$ considered collapse onto nearly the same curve, indicating that no single magnetic wave vector dominates. The system thus exhibits strong competition among various magnetic configurations, precluding the establishment of long-range magnetic order with a well-defined wave vector $\bm{Q}$.
@@ III Discussion and conclusions
The theoretical results presented in this work establish ${\mathrm{Y}}_{3}{\mathrm{Cu}}_{2}{\mathrm{Sb}}_{3}{\mathrm{O}}_{14}$ as a compelling candidate for hosting a quantum spin liquid ground state. The combination of geometric frustration inherent to the triangular lattice geometry, the presence of two inequivalent Cu sites with inverted crystal-field splittings, and the strong electronic correlations revealed by our DMFT calculations creates an ideal environment for the suppression of long-range magnetic order.
@@
Our DFT analysis demonstrates that the two distinct Cu sites exhibit fundamentally different orbital configurations. Cu-1, with its trigonal distorted coordination, follows the conventional crystal-field hierarchy with the $d_{z^{2}}$ orbital lowest in energy. In contrast, Cu-2 experiences an extreme axial compression that inverts this ordering, placing the unpaired hole in the $d_{z^{2}}$ orbital. This inversion, combined with the three-dimensional network of triangular layers, generates a complex electronic structure characterized by narrow bands at commensurate filling—a prerequisite for strong correlation effects.
@@
The DMFT calculations reveal a site-selective band renormalization at Cu-1 and Cu-2, which is deeply related to the distinct local environments of the two Cu ions. When the two ions are correctly treated in separate DMFT impurity problems, the symmetry of the twofold $E_{g}$ states at Cu-1 is preserved; these states are close to a Mott transition, while the $A_{1g}$ state at Cu-2 remains metallic across the parameter range studied. The contrasting electronic responses of the low-energy states at Cu-1 and Cu-2 may provide an explanation for the two-step spin condensation observed experimentally. The more correlated charge degrees of freedom at Cu-1 localize at a relatively higher temperature, allowing the remaining spin degrees of freedom on Cu-1 to interact and condense. In contrast, the charge at Cu-2 remains highly fluctuating until very low temperatures preventing the spins from condensing.
@@
Another significant observation is that our spin susceptibility calculations within the FLEX approximation provide direct evidence for the suppression of magnetic order. The absence of a dominant peak in $\chi_{s}(\bm{Q})$, together with the nearly $\bm{Q}$-independent behavior of the leading eigenvalues of $U_{s}\chi_{0}(\bm{Q})$, indicates that multiple magnetic instabilities compete with nearly equal strength. This competition becomes more pronounced with increasing $U$, as the variation of $\lambda_{\text{max}}$ across different wave vectors diminishes from 6% to just 1.25% when $U$ increases from 0.5 eV to 1.0 eV. Such behavior is a hallmark of a highly frustrated magnet poised at the brink of magnetic order but unable to select a unique ground state.
@@
These findings have important implications for experimental efforts to identify and characterize QSL behavior in this material. The two-step spin condensation observed experimentally [52] finds a natural explanation in our theoretical framework: the hierarchy of energy scales introduced by the two distinct Cu sites and their competing interactions may give rise to a cascade of fractionalization as temperature is lowered. Future experiments should focus on detecting the characteristic signatures of fractionalized excitations, such as spinon continua in inelastic neutron scattering, thermal conductivity plateaus, and the absence of magnetic ordering down to the lowest accessible temperatures.
In conclusion, our comprehensive theoretical investigation provides strong evidence that ${\mathrm{Y}}_{3}{\mathrm{Cu}}_{2}{\mathrm{Sb}}_{3}{\mathrm{O}}_{14}$ is a promising quantum spin liquid candidate. The interplay of distinct crystal-field environments, strong electronic correlations, and geometric frustration conspire to suppress magnetic order and stabilize an exotic ground state with fractionalized excitations. We anticipate that this material will serve as an important benchmark for understanding three-dimensional quantum spin liquids and their emergent phenomena.
@@ A.1 DFT and DMFT calculations
The first-principle calculations of ${\mathrm{Y}}_{3}{\mathrm{Cu}}_{2}{\mathrm{Sb}}_{3}{\mathrm{O}}_{14}$ were carried out in the framework of density-functional theory (DFT). We adopted the Quantum ESPRESSO (QE) implementation of DFT for the band structure calculations [63, 64]. The kinetic energy cutoff for wavefuncton and charge density was set as 80 Ry and 500 Ry, respectively. For the Brillouin-zone integrations, a Monkhorst–Pack [65] $k$-point mesh of $6\times 6\times 3$ was used. The sg15 database [66] of optimized normconserving Vanderbilt PPs (ONCV) was adopted for the non-relativistic pseudopotential.
To investigate the correlated electronic structure, we performed one-shot DFT+DMFT calculations by downfolding the DFT electronic structure to a three-band effective model using the maximally localized Wannier Functions(WFs) [67, 68] with the pw2wannier interface [53]. The DMFT impurity problem was solved using the continuous-time hybridization expansion solver [69, 70] implemented in home-made code “Package for Analyzing Correlated System ” (PACS) [54, 55] and further cross-checked with TRIQS packages [56, 57, 71]. The self-energy in real frequency was obtained by analytic continuation based on the maximum entropy [72, 73].
@@ A.2 Fluctuation exchange approximation
$H_{0}$ is the ten-band (projected by $\mathrm{Cu}$-$d$ orbitals) tight-binding model. $H_{\text{int}}$ is the simplified interaction Hamiltonian, expressed as:

### 19 Mar 2026 v2

@@ II.4 Spin Susceptibility
Figure 3: Fermi surface and spin susceptibility $\chi(\bm{q},i\omega_{m}=0)$. (a). The Fermi surface in 3D BZ and in a 2D cut through $\Gamma$ point display multiple pockets and nesting structures. (b). The spin susceptibility calculated from Eq. (2) displays multiple peaks with only small variation, which is further suppressed by electronic correlations. The blue and yellow curves correspond to different interaction parameters as shown by the legend. (c). The leading eigenvalues of $\chi_{0}U_{s}$ follow a similar structure as $\chi_{s}$ in (b), whose variation is also reduced by electronic correlations. (d). Both the leading eigenvalues of $\chi_{0}$ (the left-yaxis in red) and $\chi_{0}U_{s}$ (the right-yaxis in blue) for different momentum wave vectors fall into a single curve under the evolution of electronic correlations, with the former decreases and the latter increases with $U$.
To further shed light on the absence of long-range magnetic order in Y3Cu2Sb3O14, in this section we examine the spin susceptibility and its evolution under the influence of Coulomb interactions. The spin susceptibility, which encodes the magnetic response of the system, often serves as a key indicator of spontaneous symmetry breaking. In particular, a divergent spin susceptibility at a specific wave vector directly reveals the dominant real-space spin correlation pattern associated with that wave vector. Given the absence of long-range magnetic order down to very low temperatures in Y3Cu2Sb3O14, we anticipate a spin susceptibility that remains relatively broad and structureless across the entire BZ.
@@
From the Lehmann representation of Eq. (II.4), it is evident that the Fermi surface topology plays a crucial role in determining the divergent behavior of the spin susceptibility. When a substantial number of electrons on the Fermi surface scatter with the same wave vector, the resulting coherence can lead to a divergence of the spin susceptibility at that wave vector. Figure 3(a) displays the three-dimensional Fermi surface, with a two-dimensional cut through the $\Gamma$ point shown on the right. The Fermi surface of Y3Cu2Sb3O14 consists of several pockets of comparable size, giving rise to multiple competing nesting vectors. A few representative nesting vectors are indicated in the two-dimensional plot, each potentially contributing comparable coherence to electron scattering. As a result, the spin susceptibility exhibits a multi-peak structure with several competing maxima.
@@
The growing competition among different wave vectors in $\chi_{s}(\bm{Q})$ is further confirmed in the leading eigenvalues of $U_{s}\chi_{0}(\bm{Q})$, which appear in the denominator of Eq. (2) and drive the divergence of $\chi_{s}(\bm{Q})$ as they approach unity. Figure 3(c) shows these leading eigenvalues for the same interaction parameters used in panel (b). The variation of $\lambda_{\text{max}}(U_{s}\chi_{0}(\bm{Q}))$ across different $\bm{Q}$ is about 6% for $U=0.5$ eV, but reduces to 1.25% when $U$ is increased to 1.0 eV, despite the absolute value rising from 0.68 to 0.82. Consistent with the behavior of $\chi_{s}(\bm{Q})$, increasing interaction strength suppresses the momentum dependence of $\lambda_{\text{max}}(U_{s}\chi_{0}(\bm{Q}))$, leading to a nearly simultaneous divergence tendency at all $\bm{Q}$.
Figure 3(d) extends this analysis to more values of $U$. In addition to $U_{s}\chi_{0}(\bm{Q})$ (right y-axis, blue), we also show the leading eigenvalues of $\chi_{0}(\bm{Q})$ (see Eq. (II.4) and the left y-axis of Fig. 3(d) in red), which appear in the numerator of Eq. (2). As $U$ increases, the leading eigenvalues of $U_{s}\chi_{0}(\bm{Q})$ at characteristic wave vectors grow with nearly identical magnitudes. Concurrently, those of the bubble susceptibility $\chi_{0}(\bm{Q})$ decrease. Notably, for both quantities, the leading eigenvalues across all $\bm{Q}$ considered collapse onto nearly the same curve, indicating that no single magnetic wave vector dominates. The system thus exhibits strong competition among various magnetic configurations, precluding the establishment of long-range magnetic order with a well-defined wave vector $\bm{Q}$.
@@ III Discussion and conclusions
The theoretical results presented in this work support Y3Cu2Sb3O14 as a compelling candidate for hosting a paramagnetic ground state. The combination of geometric frustration inherent to the triangular lattice geometry, the presence of two inequivalent Cu sites with inverted crystal-field splittings, and the strong electronic correlations revealed by our DMFT calculations creates an ideal environment for the suppression of long-range magnetic order.
@@
Our DFT analysis demonstrates that the two distinct Cu sites exhibit fundamentally different orbital configurations. Cu-1, with its trigonal distorted coordination, follows the conventional crystal-field hierarchy with the $d_{z^{2}}$ orbital lowest in energy. In contrast, Cu-2 experiences an extreme axial compression that inverts this ordering, placing the unpaired hole in the $d_{z^{2}}$ orbital. This inversion, combined with the three-dimensional network of triangular layers, generates a complex electronic structure characterized by narrow bands with different filling levels.
@@
The DMFT calculations reveal a site-selective band renormalization at Cu-1 and Cu-2, which is deeply related to the distinct local environments of the two Cu ions. When the two ions are correctly treated in separate DMFT impurity problems, the symmetry of the twofold $E_{g}$ states at Cu-1 is preserved; these states are close to a Mott transition, while the $A_{1g}$ state at Cu-2 remains metallic across the parameter range studied. The contrasting electronic responses of the low-energy states at Cu-1 and Cu-2 may provide an explanation for the two-step spin condensation observed experimentally. The more correlated charge degrees of freedom at Cu-1 localize at a relatively higher temperature, allowing the remaining spin degrees of freedom on Cu-1 to interact and condense. In contrast, the charge at Cu-2 remains highly fluctuating until very low temperatures preventing the spins from condensing. The DMFT local approximation cannot fully open the charge gap. If the system is insulating from experimental point of view, a short-range magnetic correlation may play a critical role in this system.
@@
Another significant observation is that our spin susceptibility calculations within the FLEX approximation provide direct evidence for the suppression of magnetic order. The absence of a dominant peak in $\chi_{s}(\bm{Q})$, together with the nearly $\bm{Q}$-independent behavior of the leading eigenvalues of $U_{s}\chi_{0}(\bm{Q})$, indicates that multiple magnetic instabilities compete with nearly equal strength. This competition becomes more pronounced with increasing $U$ and the variation of $\lambda_{\text{max}}$ across different wave vectors is further suppressed. Such behavior is a hallmark of a highly frustrated magnet where magnetic order is unable to select a unique ground state.
@@
These findings have important implications for experimental efforts to identify and characterize QSL behavior in this material. The two-step spin condensation observed experimentally [56] finds a natural explanation in our theoretical framework: the hierarchy of energy scales introduced by the two distinct Cu sites and their competing interactions may give rise to a cascade of fractionalization as temperature is lowered. Future experiments should focus on detecting the characteristic signatures of the electronic and magnetic correlations related to the two different Cu local environment.
@@ A.1 DFT and DMFT calculations
The first-principle calculations of Y3Cu2Sb3O14 were carried out in the framework of density-functional theory (DFT). We adopted the Quantum ESPRESSO (QE) implementation of DFT for the band structure calculations [68, 69]. The kinetic energy cutoff for wavefuncton and charge density was set as 80 Ry and 500 Ry, respectively. For the Brillouin-zone integrations, a Monkhorst–Pack [70] $k$-point mesh of $9\times 9\times 9$ was used. The sg15 database [71] of optimized normconserving Vanderbilt PPs (ONCV) was adopted for the non-relativistic pseudopotential.
To investigate the correlated electronic structure, we performed one-shot DFT+DMFT calculations by downfolding the DFT electronic structure to a three-band effective model using the maximally localized Wannier Functions(WFs) [72, 73] with the pw2wannier interface [58]. The DMFT impurity problem with density-density interactions was solved using the continuous-time hybridization expansion solver [74, 75] implemented in home-made code “Package for Analyzing Correlated System ” (PACS) [59, 60] and further cross-checked with TRIQS packages [61, 62, 76]. The self-energy in real frequency was obtained by analytic continuation based on the maximum entropy [77, 78].
@@ A.2 Fluctuation exchange approximation
$H_{0}$ is the ten-band (projected by Cu-$d$ orbitals) tight-binding model. $H_{\text{int}}$ is the simplified interaction Hamiltonian, expressed as: